# Can We Verify Step by Step for Incorrect Answer Detection?

### v1

@@ Abstract
Chain-of-Thought (CoT) prompting has marked a significant advancement in enhancing the reasoning capabilities of large language models (LLMs). Previous studies have developed various extensions of CoT, which focus primarily on enhancing end-task performance. In addition, there has been research on assessing the quality of reasoning chains in CoT. This raises an intriguing question: Is it possible to predict the accuracy of LLM outputs by scrutinizing the reasoning chains they generate? To answer this research question, we introduce a benchmark, R2PE, designed specifically to explore the relationship between reasoning chains and performance in various reasoning tasks spanning five different domains. This benchmark aims to measure the falsehood of the final output of LLMs based on the reasoning steps. To make full use of information in multiple reasoning chains, we propose the process discernibility score (PDS) framework that beats the answer-checking baseline by a large margin. Concretely, this resulted in an average of $5.1\%$ increase in the F1 score across all 45 subsets within R2PE. We further demonstrate our PDS’s efficacy in advancing open-domain QA accuracy. Data and code are available at https://github.com/XinXU-USTC/R2PE.
@@
Figure 2: An overview of R2PE benchmark and the PDS framework. The construction of R2PE encompasesses 5 stages: data collection, response generation, extraction, manual cleaning, and labeling. Sourced from eight distinct datasets covering a variety of task types and derived from 6 different LLMs, R2PE comprises 45 subsets, featuring approximately 20.2K TRUE and 18.1K FALSE instances across 5 diverse domains. The objective is to establish a discernibility score that accurately indicates the veracity of the answers. PDS adopts implicit answer checking and explicit process supervision to detect all potential discrepancies among different reasoning pathways, which is superior to ADS that merely focuses on the answer consensus.
@@ 1 Introduction
Each evaluated question or claim begins with the generation of multiple rationales, leading to an answer aggregation based on majority voting (Wang et al., 2022) after extraction and manual cleaning. Once the final outcomes are labeled as true or false using an exact match, our goal is to derive a discernibility score (DS), which is intended as an indicator of the credibility of the final answer. To fully exploit the information contained in all reasoning chains, we present a process discernibility score (PDS) that substantially exceeds the answer discernibility score (ADS) baseline that counts the number of the same answers in terms of the F1 score. As shown in Figure 1, PDS can detect conflicting information about the actor and predict the incorrectness of the final answer. Figure 2 presents a succint overview of both our benchmark and the associated methodology.
@@
  • We introduce a process discernibility score (PDS) that aggregates the information in different chains of reasoning to validate the correctness of the final output.
  • Comprehensive experiments of PDS in R2PE reveal its superiority over the answer discernibility score (ADS), leading to consistent increases in the F1 score.
@@ 2.1 Construction process
Response Generation. For each question (or claim in the FEVER dataset) $Q$, the LLM is prompted using CoT to produce $n$ responses $(R_{1},R_{2},...,R_{n})$, respectively. The full prompts are given in Appendix F. The responses are aggregated from six distinct LLMs to ensure the generalizability: text-davinci-003 (Ouyang et al., 2022), GPT-3.5-turbo (OpenAI-Blog, 2022), its instruct-trained variant GPT-3.5-turbo-instruct, Gemini Pro, Mixtral-8x7b (Jiang et al., 2024), and mistral-medium. A concise exposition on LLMs along with the particulars pertaining to each model, is deferred to Appendix A.2. Specifically, we set the temperature to 0.7, $n=5$ for all datasets and ’topK’ to be 40, ’topP’ to be 0.95 in the Gemini Pro API.
Extraction. Ratinale $r_{i}$ and the corresponding anwser $a_{i}$ need to be isolated from each response $R_{i}$. During the pilot extraction period, we use the prompt cue "The answer is" as a delimiter to segregate responses. Our observations indicated that almost all LLMs yielded some outputs that did not conform to the expected answer format. Given the varying responses styles across different LLMs, the answer trigger words are identified for different LLMs and datasets after observing the original responses to facilitate extraction (see Table 6). This strategy effectively segments most of the responses.
Manual Cleaning. To maintain a high-quality benchmark, manual inspection and cleaning are performed to handle unusual cases. For atypical responses that deviate from the recognized patterns, we either manually separate the response to derive the answer or assign a special marker when separation is not feasible. Additionally, special attention has been paid to non-sentical responses. A detailed description of the extraction and manual cleaning procedure is provided in Appendix A.3.
@@ 2.2 R2PE Overview
Each instance $e$ within R2PE incorporates a series of structured data fields: question or claim $Q$, the associated dataset name, the queried LLM name, five responses alongside their reasoning chains and answers $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})\}$, the final output $a$, the ground-truth answer of question $Q$, and the label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. Detailed explications of the constituent data fields, accompanied by an illustrative example, are shown in Table 10. Additional examples are presented in Appendix B.
With R2PE, we aim to predict the correctness of final outcomes based on the intermediate reasoning steps. The response set $R$ serves as an input for predicting the label $L$ for every instance $e$. We introduce a numerical criterion, the discernibility score (DS), to encapsulate the quality of the rationales and their corresponding answers. A low DS might suggest a potential mismatch between the final output $a$ and the ground-truth answer. Hence, we will classify the example as false: $\hat{L}=F$, if its DS falls below a certain threshold $H$.
3 Process Supervision
Having R2PE at our disposal, our goal is to find a suitable DS that can be a good indicator for the veracity of final predictions. We suggest the process supervision to be taken into account recognizing its potential to overcome the limitations of DS based solely on answer agreement, which tends to overlook a great number of examples with the label $L=F$. To demonstrate its usefulness, we have conducted comprehensive experiments (results are presented in Section 4).
@@ 3.1 The Fallacy of Answer Agreement
Based on the insights of Wang et al. (2022); Zhao et al. (2023a), the degree of consensus among answers, is posited as a diagnostic tool to gauge the instances in which LLMs may make wrong predictions. We refer to the answer aggreement number as Answer discernibility Score (ADS) hereinafter. The ADS can thus be adopted as a natural baseline in our R2PE benchmark. A high ADS often suggests a high likelihood that the proposed answer is correct. Given this premise, designating a threshold of $H=n/2$ is a reasonable strategy. If the ADS is below that midpoint, $\hat{L}=F$ is predicted.
However, LLMs may generate substantially concordant but incorrect answers in reasoning tasks. Figure 1 showcases one such example, where the LLM is queried about the nationality of an actor from "Hostage for a Day". The LLM gives the answer "American" four times and "Canadian" once. Here, the ADS stands at 4, which erroneously points to a seemingly accurate but ultimately false consensus. This example serves to warn against overreliance on ADS as the sole indicator of the veracity of the final answer in reasoning tasks.
Figure 3: Qalitative analysis of results on the subset (HotpotQA from GPT-4): ADS (left), PDS (middle), and PDS-ADS (right). ADS has high precision: large answer agreement does not always guarantee accurate predictions. PDS yields desirable outcomes, while PDS-ADS has poor precision.
3.2 Process Supervision
The ADS appears to be an inadequate indicator of the ground truth label $L$, a limitation that may be ascribed to its myopic focus on the answer aggreement, while disregarding potential inconsistencies within the reasoning chain, as indicated by the areas marked red in Figure 1. Each response $R_{i}$ constitutes a rationale $r_{i}$ paried with an answer $a_{i}$. The ADS, however, limits its analysis to the answers $(a_{1},...,a_{n})$, which takes up only a small proportion of the information contained in $R$. Considering the demonstrated significance of intermediate reasoning steps (Yoran et al., 2023), a DS that makes use of rationales would be preferable.
Inspired by the hypothesis that LLMs tend to generate inaccurate or irrelevant information when they are uncertain about their predictions (Manakul et al., 2023), we introduce a process discernibility score (PDS). The PDS is designed to assess potential discrepancies among all responses, going beyond explicit verification of answers alone.
As illustrated in Figure 2, our PDS is defined as
$PDS=(\overline{ADS}+PSS)/2,$ (1)
where $\overline{ADS}=2(ADS-n/2)/n$ is a normalized ADS, which adjusts the original ADS to a balanced scale around zero. The Process Supervision Score $PSS=mean_{i\neq j}PPSS(R_{i},R_{j})$ is the average of all Pairwise PSS (PPSS). $PPSS(A,B)$ can be any score that quantifies the extent to which the content of B is included in and contradicted by A.
@@ 3.2 Process Supervision
Compared to ADS, our PDS is capable of overseeing reasoning processes as well as implicit answer checking. We would certainly expect a performance improvement over the ADS (see Section 4). As depicted in Figure 1, the PDS can detect contradictive information about the name of the actor playing a Russian hostage taker in "Hostage for a Day” among multiple reasoning chains (highlighted by red) and give a correct label $\hat{L}=F$, while the ADS only assesses the agreements among the answers and leads to a wrong prediction $\hat{L}=T$.
@@
PDS 59.64 60.68 53.01 $\textbf{63.64}^{b}$ 65.72 - +3.70
HotpotQA ADS 72.78 71.52 $63.95^{a}$ 81.46 74.24 - -
PDS 85.71 78.84 $\textbf{70.25}^{a}$ 85.65 78.66 - +7.00
2WikiMultihop ADS 69.26 69.75 $42.67^{a}$ 62.83 69.36 - -
PDS 78.65 76.51 $\textbf{57.14}^{a}$ 71.76 70.81 - +9.80
@@
Table 1: PDS consistently outperform ADS across all subsets in our R2PE benchmark. All metrics are F1 score (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively. a. These four figures are from the (HotpotQA from GPT-4) and (2WikiMultihop from GPT-4) respectively (see Section 2). b. These two are reported using a different threshold, and the details are discussed in Appendix C.
3.3 Experiment Setup
Experiments are carried out to show the effectiveness of PDS across all subsets of the R2PE benchmark. We will employ the F1 score as our evaluation metric (see Section 2), with precision and recall also included as supplementary measures. We establish the ADS as our baseline. For subsets comprising binary (yes/no) choices or multiple-choice queries, there are inherently limited answer possibilities, which often results in greater answer agreement among responses. To accommodate this characteristic, the threshold $H$ for ADS is set to 4.5 for subsets in R2PE associated with StrategyQA, Play, Physics, and FEVER. On the other hand, in subsets originated from GSM8K, MATH, HotpotQA, and 2WikiMultihop, the midpoint $H=2.5$ is chosen.
For the $PPSS(\cdot,\cdot)$ score, we adopt the SAUMMAC zero-shot model (Laban et al., 2022) that trained on MNLI (Williams et al., 2018). It is a consistency detection method for text summarization that leverages the out-of-the-blox natural language inference model to compute pairwise entailment score. As suggested by Laban et al. (2022), the probability of sentence-level entailment and contradiction is used. To detect all potential inconsistency in any sentence, we use $min$ operation to aggregate sentence-level entailment and contradiction:
$PPSS(A,B)$ $=\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})$ (2)
$\quad-con(a_{i},b_{j})\bigr{)},$
where $a_{i},b_{j}$ are sentences of $A,B$ correspondingly, $ent$ represents the probability of entailment, $con$ is the probability of contradiction, the $min$ is taken over all sentences of $B$, and $max$ is taken over all sentences of $A$. The range of this PPSS spans in $(-1,1)$. A negative value implies the potential presence of confilcting information. Consequently, it is reasonable to set the threshold $H$ for the PDS to be 0 for numeric and free form questions and $0.4$ otherwise. Comprehensive ablation studies pertaining to alternative PPSS choices are discussed in Section 4.2. The F1 score is employed to assess the performance in the various subsets of R2PE.
4 Results and Analysis
We present our findings in Table 1, which shows that PDS outperforms the baseline in all subgroups of the R2PE benchmark. To investigate the effects of the components of PDS (see Equations (1) and (2)), ablation studies are carried out in subsets related to HotpotQA and 2WikiMultihop. Finally, we demonstrate the effectiveness of our approach in enhancing the accuracy of open-domain question answering by integrating it with the verify-and-edit (VE) framework (Zhao et al., 2023a).
@@ 4 Results and Analysis
Table 2: Ablation study of PDS. All metrics are F1 score (in %). Top: subsets related to HotpotQA on R2PE. Bottom: subsets related to 2WikiMultihop on R2PE.

### v2

@@ Abstract
Chain-of-Thought (CoT) prompting has marked a significant advancement in enhancing the reasoning capabilities of large language models (LLMs). Previous studies have developed various extensions of CoT, which focus primarily on enhancing end-task performance. In addition, there has been research on assessing the quality of reasoning chains in CoT. This raises an intriguing question: Is it possible to predict the accuracy of LLM outputs by scrutinizing the reasoning chains they generate? To answer this research question, we introduce a benchmark, R2PE, designed specifically to explore the relationship between reasoning chains and performance in various reasoning tasks spanning five different domains. This benchmark aims to measure the falsehood of the final output of LLMs based on the reasoning steps. To make full use of information in multiple reasoning chains, we propose the process discernibility score (PDS) framework that beats the answer-checking baseline by a large margin. Concretely, this resulted in an average of $5.1\%$ increase in the F1 score and $2.97\%$ improvement in AUC-PR across all 45 subsets within R2PE. We further demonstrate our PDS’s efficacy in advancing open-domain QA accuracy. Data and code are available at https://github.com/XinXU-USTC/R2PE.
@@
Figure 2: An overview of R2PE benchmark and PDS framework. The construction of R2PE includes 5 stages: data collection, response generation, extraction, manual cleaning, and labeling. Sourced from 8 distinct datasets covering a variety of task types and derived from 6 different LLMs, R2PE comprises 45 subsets, featuring approximately 20.2K TRUE and 18.1K FALSE instances across 5 domains. The objective is to establish a discernibility score that accurately indicates the veracity of answers. PDS adopts answer checking and process supervision to detect all potential discrepancies among different rationales, which beats ADS that focuses merely on the answer consensus.
@@ 1 Introduction
Each evaluated question or claim begins with the generation of multiple rationales, leading to an answer aggregation based on majority voting (Wang et al., 2022) after extraction and manual cleaning. Once the final outcomes are labeled as true or false using an exact match, our goal is to derive a discernibility score (DS), which is intended as an indicator of the credibility of the final answer. To fully exploit the information contained in all reasoning chains, we present a process discernibility score (PDS) that substantially exceeds the answer discernibility score (ADS) baseline that counts the number of the same answers. As shown in Figure 1, PDS can detect conflicting information about the actor and predict the incorrectness of the final answer. Figure 2 presents a succinct overview of both our benchmark and the associated methodology.
@@
  • We introduce a process discernibility score (PDS) framework that aggregates the information in different chains of reasoning to validate the correctness of the final output.
  • Comprehensive experiments of PDS in R2PE reveal its superiority over the answer discernibility score (ADS), leading to consistent increases in the F1 score and AUC-PR.
@@ 2.1 Construction Process
Response Generation. For each question (or claim in the FEVER dataset) $Q$, the LLM is prompted using CoT to produce $n$ responses $(R_{1},R_{2},...,R_{n})$, respectively. The full prompts are given in Appendix F. The responses are aggregated from six distinct LLMs to ensure the generalizability: text-davinci-003 (Ouyang et al., 2022), GPT-3.5-turbo (OpenAI-Blog, 2022), its instruct-trained variant GPT-3.5-turbo-instruct, Gemini Pro, Mixtral-8x7b (Jiang et al., 2024), and mistral-medium. A concise exposition along with detailed settings, is deferred to Appendix A.2.
Extraction. Rationale $r_{i}$ and the corresponding anwser $a_{i}$ need to be isolated from each response $R_{i}$. During the pilot extraction period, we use the prompt cue "The answer is" as a delimiter to segregate responses. Our observations indicated that almost all LLMs yielded some outputs that did not conform to the expected answer format. Given the varying responses styles across different LLMs, the answer trigger words are identified for different LLMs and datasets after observing the original responses to facilitate extraction (see Table 7). This strategy effectively segments most of the responses.
Manual Cleaning. To maintain a high-quality benchmark, manual inspection and cleaning are performed to handle unusual cases. For atypical responses that deviate from the recognized patterns, we either manually separate the response to derive the answer or assign a special marker when separation is not feasible. Additionally, special attention has been paid to responses that are not sentences. A detailed description of the extraction and manual cleaning procedure is provided in Appendix A.3.
@@ 2.2 R2PE Overview
Each instance $e$ within R2PE incorporates a series of structured data fields: question or claim $Q$, the associated dataset name, the queried LLM name, five responses alongside their reasoning chains and answers $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})\}$, the final output $a$, the ground-truth answer of question $Q$, and the label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. Detailed explications of the constituent data fields and illustrative examples are presented in Appendix B.
With R2PE, we aim to predict the correctness of final outcomes based on the intermediate reasoning steps. The response set $R$ serves as an input for predicting the label $L$ for every instance $e$. We introduce a numerical criterion, the discernibility score (DS), to encapsulate the quality of $R$.A low DS might suggest a potential mismatch between the final output $a$ and the ground-truth answer. Hence, we will classify the example as false: $\hat{L}=F$, if its DS falls below a certain threshold $H$.
3 Process Discernibility Score
Our goal is to find a suitable DS that can be a good indicator for the veracity of final predictions. We suggest process supervision be taken into account recognizing its potential to overcome the limitations of DS based solely on answer agreement, which tends to overlook a great number of examples with the label $L=F$. To demonstrate its usefulness, we have conducted comprehensive experiments (results are presented in Section 4).
@@ 3.1 The Fallacy of Answer Agreement
Figure 3: Visualization of results on the subset (HotpotQA from GPT-4): ADS (left), PDS (middle), and PDS-ADS (right). ADS has high precision: large answer agreement does not always guarantee accurate predictions. PDS yields desirable outcomes, while PDS-ADS has poor precision.
Based on the insights of Wang et al. (2022); Zhao et al. (2023a), the degree of consensus among answers, is posited as a diagnostic tool to gauge the instances in which LLMs may make wrong predictions. We refer to the answer agreement number as Answer Discernibility Score (ADS) hereinafter. The ADS can thus be adopted as a natural baseline in our R2PE benchmark. A high ADS often suggests a high likelihood that the proposed answer is correct. Given this premise, designating a threshold of $H=n/2$ is a reasonable strategy. If the ADS is below that midpoint, $\hat{L}=F$ is predicted.
However, LLMs may generate substantially consistent but incorrect answers in reasoning tasks. Figure 1 showcases one such example, where the LLM is queried about the nationality of an actor from "Hostage for a Day". The LLM gives the answer "American" four times and "Canadian" once. Here, the ADS stands at 4, which erroneously points to a seemingly accurate but ultimately false consensus. This example serves to warn against overreliance on ADS as the sole indicator of the veracity of the final answer in reasoning tasks.
3.2 Process Discernibility Score
In our R2PE benchmark, each response $R_{i}$ constitutes a rationale $r_{i}$ paired with an answer $a_{i}$. The ADS, however, limits its analysis to the answers $(a_{1},...,a_{n})$, which takes up only a small proportion of the information contained in $R$. Considering the demonstrated significance of intermediate reasoning steps (Yoran et al., 2023), a DS that makes use of rationales would be preferable.
Inspired by the hypothesis that LLMs tend to generate inaccurate information when they are uncertain about their predictions (Manakul et al., 2023), we introduce PDS framework, which is designed to assess discrepancies among all responses, going beyond explicit verification of answers alone.
To quantify the similarity between two responses, we employ Pairwise Process Supervision Score PPSS(A, B) to measure the degree to which the content of B is both included in and opposes that of A. PPSS is confined within the interval (-1, 1), where a positive score indicates a higher level of similarity relative to contradiction. For a set of n responses, we can calculate the average of PPSS across all ordered response pairs, culminating in the Process Supervision Score (PSS): $\text{PSS}=\text{mean}_{i\neq j}\text{PPSS}(R_{i},R_{j})$. This allows PSS to evaluate the information across all responses. To underscore the final answers, we get
$\text{PDS}=(\overline{\text{ADS}}+\text{PSS})/2,$ (1)
wherein $\overline{\text{ADS}}=2(\text{ADS}-n/2)/n$ represents the normalized ADS, adjusted to align with PSS’s range. The overview of PDS is shown in Figure 2.
@@ 3.2 Process Discernibility Score
Compared to ADS, our PDS is capable of overseeing reasoning processes as well as doing implicit answer checking ($\overline{\text{ADS}}$ term in Equation (1)). As depicted in Figure 1, the PDS can detect contradictive information about the name of the actor playing a Russian hostage taker in "Hostage for a Day” among multiple reasoning chains (highlighted by red) and give a correct label $\hat{L}=F$, while the ADS only assesses the agreements among the answers and leads to a wrong prediction $\hat{L}=T$.
3.3 Experimental Setup
We will detail experimental setup in this section.
Dataset. All subsets of the R2PE benchmark.
Baseline. From Wang et al. (2022); Zhao et al. (2023a), ADS can act as the baseline in our setting.
Metrics. We will employ the F1 score and AUC-PR as our evaluation metrics, with precision and recall also included as supplementary measures. As we aim to detect potential incorrect answers, False labels are treated as positive. Therefore, true positives are examples with $L=F$ and $\hat{L}=F$.
PDS Implementation. For $\text{PPSS}(\cdot,\cdot)$, we adopt the SAUMMAC zero-shot model (Laban et al., 2022) that trained on MNLI (Williams et al., 2018), which is a consistency detection method for text summarization that leverages the out-of-the-box natural language inference model to compute pairwise entailment score. As suggested by Laban et al. (2022), probabilities of sentence-level entailment and contradiction are used. To detect all potential inconsistencies in any sentence, we use $min$ operation to aggregate across sentence pairs:
$\text{PPSS}(A,B)$ $=\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})$ (2)
$\quad-con(a_{i},b_{j})\bigr{)},$
where $a_{i},b_{j}$ are sentences of $A,B$ correspondingly, $ent$ represents the probability of entailment, $con$ is the probability of contradiction, the $min$ is taken over all sentences of $B$, and $max$ is taken over all sentences of $A$. Note that the PPSS metric lacks symmetry; however, this is not a problem in our implementation because the PSS calculation averages across all possible ordered response pairs.
Threshold Selection. As $\text{ADS}\in\{0,1,2,...,5\}$, threshold $H$ is set to be the midpoint $H=2.5$. Another reason for this selection is that $\text{ADS}>2.5$ means the majority agrees, which aligns the choice in Zhao et al. (2023a). The range of this PPSS spans in $(-1,1)$. A negative value implies the potential presence of conflicting information. Consequently, it is reasonable to set the threshold $H$ for the PDS to be 0. Our choice of threshold $H$ is not task-specific. We regard this as a benefit because held-out data for threshold tuning may not always be available in the real scenario. A detailed discussion is given in Appendix C.3.
@@ 3.3 Experimental Setup
PDS 59.64 60.68 53.01 63.64 65.72 - +3.70
HotpotQA ADS 72.78 71.52 63.95 81.46 74.24 - -
PDS 85.71 78.84 70.25 85.65 78.66 - +7.00
2WikiMultihop ADS 69.26 69.75 42.67 62.83 69.36 - -
PDS 78.65 76.51 57.14 71.76 70.81 - +9.80
@@
Table 1: PDS consistently outperforms ADS across all subsets in our R2PE benchmark in F1 scores (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively.
Dataset Method GPT3 GPT-instruct GPT-3.5 Gemini Mixtral mistral avg
GSM8K ADS 70.79 55.83 54.58 52.59 63.18 48.15 -
PDS 74.14 61.02 55.37 56.23 68.37 51.86 +3.66
MATH ADS 92.82 85.50 78.28 87.81 90.13 88.05 -
PDS 93.83 86.77 80.84 89.49 91.82 89.66 +1.64
StrategyQA ADS 49.56 51.42 49.36 49.85 59.55 56.27 -
PDS 49.97 54.49 52.87 51.18 62.08 67.22 +3.63
Play ADS 44.37 48.93 49.90 46.00 72.68 50.44 -
PDS 46.14 48.01 51.31 47.66 72.84 51.86 +0.92
Physics ADS 39.94 40.60 48.63 79.10 53.05 52.58 -
PDS 46.85 48.58 61.31 78.50 56.99 60.66 +6.50
FEVER ADS 57.02 49.72 48.57 54.02 67.31 - -
PDS 62.19 51.41 49.00 55.43 66.91 - +1.66
HotpotQA ADS 89.71 79.14 83.99 90.15 74.89 - -
PDS 92.61 83.21 86.21 91.47 79.41 - +3.01
2WikiMultihop ADS 83.26 75.52 48.12 79.37 59.51 - -
PDS 86.72 79.80 54.37 80.20 58.94 - +2.38
avg - +3.12 +3.33 +3.74 +1.41 +1.91 +5.15 +2.97
Table 2: PDS consistently outperforms ADS across almost all subsets in our R2PE benchmark in AUC-PR (in %).
@@
Table 3: Ablation study of PDS. All metrics are F1 scores (in %). Top: subsets related to HotpotQA on R2PE. Bottom: subsets related to 2WikiMultihop on R2PE.
4 Results and Analysis